# DecoR: Deconfounding Time Series with Robust Regression

### v1

@@ D.1 Lemmas
Since |$|\{S\subseteq\{1,\dots,n\}\text{ s.t. }|S|=a_{n}\}|=\binom{n}{a_{n}}=\binom{n}{c_{n}}\leq(en/c_{n})^{c_{n}}$ and by Lemma D.1 and the union bound (see the proof of Theorem 3.1 for a similar argument), there exists $K>0$ such that with probability at least $1-\delta$ for all $t\in\mathbb{N}$ that
@@
By Theorem 3.3 and the fact that $\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}=\left(\lambda_{\mathrm{min}}\left(X_{S_{t}}^{T}X_{S_{t}}\right)\right)^{-1/2}$ we have
@@
This yields the first result. If the rows of $X$ are i.i.d. standard Gaussian random vectors, we have $\lambda_{a_{n}}(X)\in\Omega(\sqrt{n})$ (see Bhatia et al. (2015, Theorem 15)). This yields the second statement.
D.6 Proof of Theorem 4.1 and Theorem 4.2
Lemma D.4.
Assume the setup of Section 4, let $c_{n}$ be a sequence satisfying, for all $n$, $|G_{n}|\leq c_{n}$ and define the $\mathcal{U}_{n}$ as in Theorem 4.1 and let Assumption 2 be satisfied. Then, the following three statements hold.
@@
    for all $n\in\mathbb{N}$ the $n$ components of $\eta^{n}_{\phi}$ are i.i.d. centered Gaussian random variables with variance $\bar{\sigma}^{2}=\sigma_{\eta}^{2}/n$,
@@
Since $\{T_{k}^{\phi,n}(\eta)\}_{k\leq n}$ are linear combinations of independent centered Gaussians, they are jointly Gaussian with mean zero. By Assumption 2 (i) it holds that
@@
Proof of Theorem 4.1.
We want to apply Theorem 3.1. Since for all $n\in\mathbb{N}$ and $j,i\leq T$ it holds that $\frac{1}{n}\sum_{l=1}^{n}\phi_{l}(Tj/n)\phi_{l}(Ti/n)=\mathbbm{1}{\{i=j\}}$ (by assumption) we have that
@@
By (24) and Lemma D.4 (iii) for all $\delta>0$ there exist $v_{1},v_{2}>0$ and $\bar{n}$ such that or all $n\geq\bar{n}$, with probability at least $1-\delta$, for all $S\in\mathcal{U}_{n}$, we have $||(X^{n}_{\phi})_{S\setminus G_{n}}||_{2}>v_{1}$ and $||(X^{n}_{\phi})_{S}||_{2}<v_{2}$. By definition of $\mathcal{U}_{n}$ we have that for all $S\in\mathcal{U}_{n}$ and all $U_{n}\in\mathrm{Inl}(\mathcal{U}_{n})$ it holds that $|S|=|U_{n}|=n-c_{n}$ and therefore $\alpha_{1}(S,U_{n},\delta)\leq 4\sigma^{2}\sqrt{n}\sqrt{2\log(2|\mathcal{U}_{n}|/\delta)}$. Property (i) and Property (ii) of Lemma D.4 then yield the claim by Theorem 3.1. ∎
Proof of Theorem 4.2.
@@
and by (iii) we know that for large $n$ we have with high probability that for all for all $S$ with $|S|=n-c_{n}$ the value of $\sqrt{\lambda_{\mathrm{min}}\left((X^{n}_{\phi})_{S}^{T}(X^{n}_{\phi})_{S}\right)}$ is lower bounded. This yields the desired result by Corollary 3.4. ∎
D.7 Proof of Proposition 4.3
@@
By Lemma D.4 (ii) the first term vanishes in probability. The second term, however, is non-vanishing in many cases, for example, when $U$ is a band-limited process and $X_{t}=U_{t}+\hat{\epsilon}_{t}$ for all $t\in[0,T]$, where $\hat{\epsilon}$ is a band-limited process independent of $U$.
@@ Appendix E Additional Results
E.2 Example application of Theorem 3.1
If (7) holds, all terms in the bound of the estimation error in Theorem 3.1 are asymptotically bounded by $\sqrt{\log(\mathcal{U}_{n})/\delta}/\sqrt{n}$, except the term that includes $\alpha_{1}$. If we multiply out the quadratic terms appearing in $\alpha_{1}$, we obtain
@@ E.3 Lower Bound
Assume we observe $\{(x_{i},\tilde{Y}_{i})\}_{i\in\mathbb{N}}$, where $\tilde{Y}_{1},\ldots,\tilde{Y}_{n}$ are obtained by (adversarially) perturbing $c_{n}$ of the $Y$-values. If there exists $c,o\in(0,1)$ and $\bar{n}\in\mathbb{N}$ such that for all $n\geq\bar{n}$
@@
Assume that $X_{1},X_{2},\dots$ are i.i.d. distributed such that for all $i\in\mathbb{N}$ we have $\mathbb{P}[X_{i}=1]=\mathbb{P}[X_{i}=-1]=1/2$ and let $c_{n}=\lfloor\eta n\rfloor$ for $\eta\in(0,1]$ (that is, $\eta$ is the fraction of outliers). Then $p_{n}\leq n|\beta_{1}-\beta_{2}|$ and for $|\beta_{1}-\beta_{2}|$ small enough there exists $o>0$ such that for all $n$, $p_{n}/\lfloor\eta n\rfloor<1-o$. Therefore, by Theorem E.1(in general) there does not exist a consistent estimator if we allow for a constant fraction of outliers.
@@ F.1 Experimental Details
For the synthetic experiments, we generate data from the model specified in Setting 1 with $\beta=3$ and $T=1$. We sample $\epsilon_{X}$ and $U$ either from two Ornstein-Uhlenbeck processes with parameters $(1,-0.8)$ and $(1,-0.5)$, respectively, or from two band-limited processes with coefficients drawn i.i.d. standard Gaussian with $S=\{1,\dots,50\}$. We choose the confounded components $G$ uniformly at random with probability $0.25$. For the othonormal basis $\phi$ we consider the cosine basis (see Definition C.1) and the Haar basis (see Definition C.2). We choose the threshold parameter $a$ for Torrent and BFS as $0.7$. We consider noise with variances $\sigma_{\eta}^{2}\in\{0,1,2\}$. We evaluate the mean absolute prediction error on 1000 sample sets, that is, $\mathrm{MAE}=\sum_{i=1}^{1000}|\beta-\hat{\beta}^{1000}_{i}|/{1000}$. In all experiments, except for the one presented in Table 1, we use DecoR-Tor.

### v2

@@ D.5 Proof of Corollary 1
Since |$|\{S\subseteq\{1,\dots,n\}\text{ s.t. }|S|=a_{n}\}|=\binom{n}{a_{n}}=\binom{n}{c_{n}}\leq(en/c_{n})^{c_{n}}$ and by Lemma 2 and the union bound (see the proof of Theorem 1 for a similar argument), there exists $K>0$ such that with probability at least $1-\delta$ for all $t\in\mathbb{N}$ that
@@
By Theorem 2 and the fact that $\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\right\rVert_{2}=\left(\lambda_{\mathrm{min}}\left(X_{S_{t}}^{T}X_{S_{t}}\right)\right)^{-1/2}$ we have
@@
This yields the first result. If the rows of $X$ are i.i.d. standard Gaussian random vectors, we have $\lambda_{a_{n}}(X)\in\Omega_{\mathbb{P}}(\sqrt{n})$ (see Bhatia et al. (2015, Theorem 15)). This yields the second statement.
D.6 Proof of Theorem 3 and Theorem 4
Lemma 4.
Assume the setup of Section 4, let $c_{n}$ be a sequence satisfying, for all $n$, $|G_{n}|\leq c_{n}$ and define the $\mathcal{U}_{n}$ as in Theorem 3 and let Assumption 2 be satisfied. Then, the following three statements hold.
@@ D.6 Proof of Theorem 3 and Theorem 4
    for all $n\in\mathbb{N}$ the $n$ components of $\eta^{n}_{\phi}$ are i.i.d. centred Gaussian random variables with variance $\bar{\sigma}^{2}=\sigma_{\eta}^{2}/n$,
@@
Since $\{T_{k}^{\phi,n}(\eta)\}_{k\leq n}$ are linear combinations of independent centred Gaussians, they are jointly Gaussian with mean zero. By Assumption 2 (i) it holds that
@@
Proof of Theorem 3.
We want to apply Theorem 1. Since for all $n\in\mathbb{N}$ and $j,i\leq T$ it holds that $\frac{1}{n}\sum_{l=1}^{n}\phi_{l}(Tj/n)\phi_{l}(Ti/n)=\mathbbm{1}{\{i=j\}}$ (by assumption) we have that
@@
By (27) and Lemma 4 (iii) for all $\delta>0$ there exist $v_{1},v_{2}>0$ and $\bar{n}$ such that or all $n\geq\bar{n}$, with probability at least $1-\delta$, for all $S\in\mathcal{U}_{n}$, we have $||(X^{n}_{\phi})_{S\setminus G_{n}}||_{2}>v_{1}$ and $||(X^{n}_{\phi})_{S}||_{2}<v_{2}$. By definition of $\mathcal{U}_{n}$ we have that for all $S\in\mathcal{U}_{n}$ and all $U_{n}\in\mathrm{Inl}(\mathcal{U}_{n})$ it holds that $|S|=|U_{n}|=n-c_{n}$ and therefore $\alpha_{1}(S,U_{n},\delta)\leq 4\sigma^{2}\sqrt{n}\sqrt{2\log(2|\mathcal{U}_{n}|/\delta)}$. Property (i) and Property (ii) of Lemma 4 then yield the claim by Theorem 1. ∎
Proof of Theorem 4.
@@
and by Lemma 4 (iii) (using the fact that $\lambda_{\mathrm{min}}$ is superadditive for positive semi-definite matrices) we know that for large $n$ we have with high probability that for all $S$ with $|S|=n-c_{n}$ the value of $\sqrt{\lambda_{\mathrm{min}}\left((X^{n}_{\phi})_{S}^{T}(X^{n}_{\phi})_{S}\right)}$ is lower bounded. This yields the desired result by Corollary 1. ∎
D.7 Proof of Proposition 1
@@
By Lemma 4 (ii) the first term vanishes in probability. The second term, however, is non-vanishing in many cases, for example, when $U$ is a band-limited process and $X_{t}=U_{t}+\hat{\epsilon}_{t}$ for all $t\in[0,T]$, where $\hat{\epsilon}$ is a band-limited process independent of $U$.
@@ Appendix E Additional Results
E.2 Example application of Theorem 1
If (7) holds, all terms in the bound of the estimation error in Theorem 1 are asymptotically bounded by $\sqrt{\log(\mathcal{U}_{n})/\delta}/\sqrt{n}$, except the term that includes $\alpha_{1}$. If we multiply out the quadratic terms appearing in $\alpha_{1}$, we obtain
@@ E.3 Lower Bound
Assume we observe $\{(x_{i},\tilde{Y}_{i})\}_{i\in\mathbb{N}}$, where $\tilde{Y}_{1},\ldots,\tilde{Y}_{n}$ are obtained by (adversarially) perturbing $c_{n}$ of the $Y$-values. if there exist $c,o\in(0,1)$ and $\bar{n}\in\mathbb{N}$ such that for all $n\geq\bar{n}$
@@
Assume that $X_{1},X_{2},\dots$ are i.i.d. distributed such that for all $i\in\mathbb{N}$ we have $\mathbb{P}[X_{i}=1]=\mathbb{P}[X_{i}=-1]=1/2$ and let $c_{n}=\lfloor\eta n\rfloor$ for $\eta\in(0,1]$ (that is, $\eta$ is the fraction of outliers). Then $p_{n}\leq n|\beta_{1}-\beta_{2}|$ and for $|\beta_{1}-\beta_{2}|$ small enough there exists $o>0$ such that for all $n$, $p_{n}/\lfloor\eta n\rfloor<1-o$. Therefore, by Theorem 6(in general) there does not exist a consistent estimator if we allow for a constant fraction of outliers.
@@ F.1 Experimental Details
For the synthetic experiments, we generate data from the model specified in Setting 1 with $\beta=3$ and $T=1$. We sample $\epsilon_{X}$ and $U$ either from two Ornstein-Uhlenbeck processes with parameters $(1,-0.8)$ and $(1,-0.5)$, respectively, or from two band-limited processes with coefficients drawn i.i.d. standard Gaussian with $S=\{1,\dots,50\}$. We choose the confounded components $G$ uniformly at random with probability $0.25$. For the othonormal basis $\phi$ we consider the cosine basis (see Definition 2) and the Haar basis (see Definition 3). We choose the threshold parameter $a$ for Torrent and BFS as $0.7$. We consider noise with variances $\sigma_{\eta}^{2}\in\{0,1,4\}$. We evaluate the mean absolute prediction error on 1000 sample sets, that is, $\mathrm{MAE}=\sum_{i=1}^{1000}|\beta-\hat{\beta}^{1000}_{i}|/{1000}$. In all experiments, except for the one presented in Table 1, we use DecoR-Tor.